mathslover (mathslover):

A particle moves in a straight line obeying the relation x = t(t-1) where x - displacement in m and t = time in sec. What are the distance covered and displacement during t = 2 sec?

mathslover (mathslover):

I solved for displacement just by putting t = 2 sec. in the given relation as x is the displacement itself. So, x = 2(2-1) = 2 m.

mathslover (mathslover):

But, not sure for the distance. @Hero @AravindG @zepdrix @austinL @mathmale

OpenStudy (shadowlegendx):

Wish I could help @mathslover :/ Good Luck with finding the answer!

mathslover (mathslover):

np Shadow. Thanks!

OpenStudy (mathmale):

May I assume that you're taking Calculus for at least have some experience with the derivative as a rate of change / velocity?

mathslover (mathslover):

Well, yes, I know basics of calculus. And yes, I have a lot of experience with differentiation but not a great experience with integration.

mathslover (mathslover):

Though, as per this question/topic is concerned, my calculus concepts are at right point of my mind Sir.

OpenStudy (mathmale):

You're asked to find the "distance covered" over (I'm assuming) the time interval [0,2]. Note that the particle can move both to the left and to the right from its starting point on the x-axis. Although its displacement from t=0 to t=2 can be found by subtracting t(0) from t(2), its total distance traveled cannot. How are you going to find the distance traveled in the (+) direction separately from the distance traveled in the negative direction? You must do so, and then must add those two distances together. So: differentiate x=t(t-1); set the derivative = to 0 and solve for t (your critical value). Why? What next?

mathslover (mathslover):

That is what confused me. If I differentiate x = t(t-1) => \(\dfrac{d(t^2 - t)}{dt} = 2t - 1\) If I set the derivative = 0 , so, I get 2t - 1 = 0 and t = 1/2

mathslover (mathslover):

I'm not sure that what does it signify? Does it signify that velocity of the particle will be zero at t= 1/2 ?

OpenStudy (mathmale):

Yes. For a clearer picture of what is happening, treat x=1/2 as a "critical value:" graph it on a number line subdivided into intervals [0,1/2) U (1/2,2] Determine for which time interval the velocity is negative and for which it is positive. How far does the particle move during the time interval [0,1/2)? during the time interval (1/2,2]? Add together the absolute values of those two distances to obtain the total distance traveled.

OpenStudy (mathmale):

Unfortunately, I need to get off the 'Net. But I'll be back on later today and could possibly help you again later if need be.

mathslover (mathslover):

Sure Sir. Thanks a lot, I will try to solve it from here.

OpenStudy (anonymous):

May i suggest u go through this video? xD http://www.youtube.com/watch?v=8Ig-e509uRQ right at 27:41.. see what he does.. its really great!!.. just go through it once :)

mathslover (mathslover):

Thanks a lot @Mashy for sharing the link. I will have a look at it.

OpenStudy (anonymous):

sure.. and do tell me what answer u get.. ll check.. if its right .. also try drawing xt, and vt graphs for this.. (to get a better insight ).. hope u see xt would be a parabola, vt would be a straight line and a is constant hence horizontal line!

mathslover (mathslover):

Yep. I drew the xt graph and I got it as a parabola ... the max. negative value for the distance being at time = 1/2 . And yes, vt graph will be a straight line. I got 2.5 m as the answer, is it right?

OpenStudy (anonymous):

yes .. at t=0 x =0 at t = 0.5 x = -0.25 hence it traveled 0.25 meters in negative x at t = 2, x = 2 hence it traveled back 0.25 and then traveled additional 2 meters both in positive x so together its 2+0.25+0.25 = 2.5 m!

OpenStudy (anonymous):

But.. ur question wording "during t = 2 sec? " it makes me think.. that it needs u to find the distances and displacement DURING the 2nd second.. from t =2s to t = 3s .. why is it worded like that :-/

OpenStudy (anonymous):

" t =2s to t = 3s ." i meant " t =1s to t = 2s ".

mathslover (mathslover):

Not sure, but the book's answer is also 2.5 , so, I guess it meant 0 to 2 sec.

OpenStudy (anonymous):

oh ok ok :)

OpenStudy (mathmale):

Again: Please be certain to distinguish between 1) total distance traveled, regardless of direction of travel, and 2) the displacement over the period [0,2]. "the book's answer" ... for what? is 2.5. This problem calls for two separate answers, as described above.

mathslover (mathslover):

Sir, 2.5 for distance travelled (total)

mathslover (mathslover):

And sorry for not being specific.